OpenStudy (anonymous):

A bag contains 9 red marbles, 10 white, 5 blue. If 6 are drawn at random, what is the probability a) 3 blue b) same number of each colour c) NONE WILL BE RED d) all 6 will be the same colour

OpenStudy (anonymous):

a.\[C{5}^{3}/C{24}^{6}\]

OpenStudy (anonymous):

two ways to do first one. \[\frac{\dbinom{5}{3}\times \dbinom{19}{3}}{\dbinom{24}{6}}\] probably easiest. but cancel your brains out before computing

OpenStudy (anonymous):

c is easiest. none will be red is just \[\frac{15}{24}\times \frac{14}{23}\times \frac{13}{22}\times \frac{12}{21}\times \frac{11}{20}\times \frac{10}{19}\]

OpenStudy (anonymous):

same color of each number means 2 of each. so number 2 is \[\frac{\dbinom{9}{2}\times \dbinom{10}{2}\times \dbinom{5}{2}}{\dbinom{24}{6}}\]

OpenStudy (anonymous):

hope the method is clear and also hope it is clear how to compute these numbers

OpenStudy (anonymous):

factorials should lead to much cancellation, so cancel first multiply last